# Reducing Events to Augment Log-based Anomaly Detection Models: An Empirical Study

### v1

@@ 2. Introduction
Modern software systems are becoming increasingly complex, leading to more frequent failures that can cause considerable losses even during short periods of unavailability(Elliot, 2014; Zhang et al., 2024b). Detecting anomalies accurately has therefore become critical for ensuring reliable and continuously available services. System logs provide valuable runtime information about software states and events, making them an indispensable resource for log-based anomaly detection approaches. With the ability to pinpoint failures and prevent further deterioration, log-based anomaly detection have garnered significant attention as important ways to maintain highly secure and resilient software systems in the face of rising complexity.
In recent years, anomaly detection based on system logs has gained significant research attention. These log-based anomaly detection models can be broadly classified into two categories: supervised models(Chen et al., 2004; Fronza et al., 2013; Bodik et al., 2010; Liang et al., 2007; Zhang et al., 2019; Yang et al., 2021; Lu et al., 2018; Zhang et al., 2024a) and unsupervised models(Babenko et al., 2009; Du et al., 2017; Yin et al., 2020; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022; Kim et al., 2020; Meng et al., 2019; Landauer et al., 2023). Supervised models, such as RobustLog(Zhang et al., 2019), necessitate labeled data comprising both normal and abnormal instances to construct their predictive frameworks. In contrast, unsupervised models detect deviations relying solely on standard data. They are primarily split into deep neural network-based(Du et al., 2017; Kim et al., 2020; Meng et al., 2019; Yin et al., 2020) and graph-based models(Babenko et al., 2009; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022; Kim et al., 2020).
@@
To fill this significant gap, we conduct an empirical study to quantify the impact of log reduction on anomaly detection The investigation spans six anomaly detection models (LR(Bodik et al., 2010), SVM(Liang et al., 2007), Decision Tree(Chen et al., 2004), Isolation Forest(Hariri et al., 2019), RobustLog(Zhang et al., 2019), PLELog(Yang et al., 2021)) applied to three datasets(He et al., 2023) (HDFS, BGL, Thunderbird). We design two approaches: a retry-based method and a clustering-based approach, to validate the extent of log event reduction possible under constrained model performance degradation thresholds. The results reveal that for anomaly detection models, log events can be significantly reduced, and the reduction of logs can even enhance model effectiveness. In extreme cases, such as the Thunderbird dataset, a single log event (originally 1406 log events) can identify most anomalies.
@@ 3.1. Anomaly Detection
Anomaly detection(Chen et al., 2004; Fronza et al., 2013; Bodik et al., 2010; Liang et al., 2007; Zhang et al., 2019; Yang et al., 2021; Lu et al., 2018; Babenko et al., 2009; Du et al., 2017; Yin et al., 2020; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022) aims to identify irregularities in system behavior. Detection methods are broadly categorized into supervised and unsupervised models. Supervised models(Chen et al., 2004; Fronza et al., 2013; Bodik et al., 2010; Liang et al., 2007; Zhang et al., 2019; Yang et al., 2021; Lu et al., 2018), like RobustLog(Zhang et al., 2019), require labeled data that includes both normal and abnormal examples to form predictive frameworks. PLELog(Yang et al., 2021) addresses the issue of insufficient labels via probabilistic label estimation and designs an attention-based GRU neural network to detect anomalies. Loglizer(He et al., 2016a) offers a comprehensive toolkit featuring several machine-learning-based log analysis models designed for automated anomaly detection, including linear regression (LR)(Bodik et al., 2010), SVM(Liang et al., 2007), Decision Tree(Chen et al., 2004), Isolation Forest(Hariri et al., 2019). In contrast, unsupervised models(Babenko et al., 2009; Du et al., 2017; Yin et al., 2020; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022) identify deviations based only on standard data. In this paper, we focus on supervised models in analysis, given that labeled data allows a systematic assessment of log event reduction’s impact on model performance. Unsupervised approaches cannot conclusively link performance changes to specific log event removals due to the absence of labels.
@@ 5.1. RQ1: Reduction Extent of Log Events for Anomaly Detection Methods
Subsequently, we conduct further experiments to investigate how the quantity of reduced logs varies with changing values of $\alpha$. As depicted in figure 5, for the HDFS dataset, the amount of reduction doesn’t significantly change when $\alpha\in\{0,0.01\}$. Beyond this range, it remains relatively stable. This is attributed to the fact that when $\alpha$ is too low, the results might be swayed by the model’s inherent random fluctuations, making it challenging to accurately determine the relevance of each log event.
For the BGL dataset, a similar pattern emerges, with exceptions observed for RobustLog and Isolation Forest. In the case of RobustLog, a noticeable change occurs when $\alpha=0.05$. This suggests that there’s a set of events that can be eliminated for RobustLog, but doing so might genuinely impact its performance. As for the Isolation Forest, its behavior remains consistent throughout, likely because it struggles to effectively identify anomalies in the BGL dataset.
@@
5.2. RQ2: Performance Following Event Reduction
@@ 8.1. Log-based Anomaly Detection
Log analysis for anomaly detection is a well-established research area(Du et al., 2017; Jia et al., 2021, 2022; Zhang et al., 2019; Meng et al., 2019; Duan et al., 2023; Yang et al., 2021; Chen et al., 2020; Han and Yuan, 2021; Liu et al., 2022). These methodologies typically involve extracting templates and key information from logs, followed by constructing models for anomaly detection and classification. There are mainly two types of models in this domain: graph-based and deep-learning models.
@@ 8.2. Log Compression & Placement
Given the substantial volume of logs generated by modern systems, assisting developers in adding appropriate logging statements is a promising research area, as highlighted in prior studies(Yuan et al., 2012a, b; Zhao et al., 2017; Yu et al., 2023a). Errlog(Yuan et al., 2012a), LogEnhancer(Yuan et al., 2012b), and Log20(Zhao et al., 2017) enhance debugging capabilities by strategically inserting supplementary logging statements into the source code. Concurrently, LogReducer(Yu et al., 2023a) leverages eBPF to manage logging overhead in performance-critical areas, ensuring that logging remains effective.
However, the process of archiving massive volumes of logs over extended periods can introduce substantial storage overhead. To address this challenge, several studies have focused on log compression techniques to reduce storage requirements. Approaches such as Nanolog(Yang et al., 2018), CLP(Rodrigues et al., 2021), and Cowic(Lin et al., 2015) construct dictionaries for fields in logs and replace strings by referencing these dictionaries. Additionally, LogZip(Liu et al., 2019) and RoughLogs(Meinig et al., 2019) employ sophisticated statistical models to identify and reduce redundancy in logs.
@@ 9.1. Application of LogCleaner
LogCleaner has been implemented for a subset of users in Apache IoTDB, yielding positive feedback. Beyond its application as described in this paper, some users have employed LogCleaner to identify key events and alert developers about unnecessary print statements in the logs that can be removed. Developers can selectively delete these prints to enhance system performance. It has aided developers in discovering that over 50% of the log print statements in the system are unnecessary. As a result, the performance of Apache IoTDB has improved by approximately 8%.

### v2

@@ 2. Introduction
Modern software systems are becoming increasingly complex, leading to more frequent failures that can cause considerable losses even during short periods of unavailability(Elliot, 2014; Zhang et al., 2024c). Detecting anomalies accurately has therefore become critical for ensuring reliable and continuously available services. System logs provide valuable runtime information about software states and events, making them an indispensable resource for log-based anomaly detection approaches. With the ability to pinpoint failures and prevent further deterioration, log-based anomaly detection have garnered significant attention as important ways to maintain highly secure and resilient software systems in the face of rising complexity.
In recent years, anomaly detection based on system logs has gained significant research attention. These log-based anomaly detection models can be broadly classified into two categories: supervised models(Zhang et al., 2019; Yang et al., 2021; Lu et al., 2018; Zhang et al., 2024b) and unsupervised models(Babenko et al., 2009; Du et al., 2017; Yin et al., 2020; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022; Kim et al., 2020; Meng et al., 2019; Landauer et al., 2023). Supervised models, such as RobustLog(Zhang et al., 2019), necessitate labeled data comprising both normal and abnormal instances to construct their predictive frameworks. In contrast, unsupervised models detect deviations relying solely on standard data. They are primarily split into deep neural network-based(Du et al., 2017; Kim et al., 2020; Meng et al., 2019; Yin et al., 2020) and graph-based models(Babenko et al., 2009; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022; Kim et al., 2020).
@@
To fill this significant gap, we conduct an empirical study to quantify the impact of log reduction on anomaly detection The investigation spans six anomaly detection models (LR, SVM, Decision Tree, Isolation Forest, RobustLog(Zhang et al., 2019), PLELog(Yang et al., 2021)) applied to three datasets(He et al., 2023) (HDFS, BGL, Thunderbird). We design two approaches: a retry-based method and a clustering-based approach, to validate the extent of log event reduction possible under constrained model performance degradation thresholds. The results reveal that for anomaly detection models, log events can be significantly reduced, and the reduction of logs can even enhance model effectiveness. In extreme cases, such as the Thunderbird dataset, a single log event (originally 1406 log events) can identify most anomalies.
@@ 3.1. Anomaly Detection
Anomaly detection(Zhang et al., 2019; Yang et al., 2021; Lu et al., 2018; Babenko et al., 2009; Du et al., 2017; Yin et al., 2020; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022) aims to identify irregularities in system behavior. Detection methods are broadly categorized into supervised and unsupervised models. Supervised models(Zhang et al., 2019; Yang et al., 2021; Lu et al., 2018), like RobustLog(Zhang et al., 2019), require labeled data that includes both normal and abnormal examples to form predictive frameworks. PLELog(Yang et al., 2021) addresses the issue of insufficient labels via probabilistic label estimation and designs an attention-based GRU neural network to detect anomalies. Loglizer(He et al., 2016a) offers a comprehensive toolkit featuring several machine-learning-based log analysis models designed for automated anomaly detection, including linear regression (LR), SVM, Decision Tree, Isolation Forest. In contrast, unsupervised models(Du et al., 2017; Yin et al., 2020; Jia et al., 2021; Guo et al., 2021; Jia et al., 2022) identify deviations based only on standard data. In this paper, we focus on supervised models in analysis, given that labeled data allows a systematic assessment of log event reduction’s impact on model performance. Unsupervised approaches cannot conclusively link performance changes to specific log event removals due to the absence of labels.
@@ 5.1. RQ1: Reduction Extent of Log Events for Anomaly Detection Methods
We further investigate how the quantity of reduced logs changes with varying values of $\alpha$. As shown in Figure 5, for the HDFS dataset, the reduction remains stable when $\alpha\in{0,0.01}$ and beyond, as lower values of $\alpha$ may be influenced by the model’s random fluctuations, making it hard to assess log event relevance accurately.
A similar trend is observed for the BGL dataset, with RobustLog showing a significant change at $\alpha=0.05$, indicating that some events can be removed, but this may affect performance. In contrast, Isolation Forest remains consistent, likely due to its difficulty in detecting anomalies in the BGL dataset.
@@
5.2. RQ2: Performance After Event Reduction
@@ 8.1. Log-based Anomaly Detection
Log analysis for anomaly detection is a well-established research area(Du et al., 2017; Jia et al., 2021, 2022; Zhang et al., 2019; Meng et al., 2019; Yang et al., 2021; Han and Yuan, 2021; Liu et al., 2022). These methodologies typically involve extracting templates and key information from logs, followed by constructing models for anomaly detection and classification. There are mainly two types of models in this domain: graph-based and deep-learning models.
@@ 8.2. Log Compression & Placement
Given the substantial volume of logs generated by modern systems, assisting developers in adding appropriate logging statements is a promising research area, as highlighted in prior studies(Zhao et al., 2017; Yu et al., 2023a). Log20(Zhao et al., 2017) enhance debugging capabilities by strategically inserting supplementary logging statements into the source code. Concurrently, LogReducer(Yu et al., 2023a) leverages eBPF to manage logging overhead in performance-critical areas, ensuring that logging remains effective.
However, the process of archiving massive volumes of logs over extended periods can introduce substantial storage overhead. To address this challenge, several studies have focused on log compression techniques to reduce storage requirements. Approaches such as Nanolog(Yang et al., 2018) and CLP(Rodrigues et al., 2021) construct dictionaries for fields in logs and replace strings by referencing these dictionaries. Additionally, LogZip(Liu et al., 2019) and RoughLogs(Meinig et al., 2019) employ sophisticated statistical models to identify and reduce redundancy in logs.
@@ 9.1. Application of LogCleaner
LogCleaner has been implemented for a subset of users in Apache IoTDB (ZHANG et al., 2021; Kang et al., 2022; Zhang et al., 2024a), yielding positive feedback. Beyond its application as described in this paper, some users have employed LogCleaner to identify key events and alert developers about unnecessary print statements in the logs that can be removed. Developers can selectively delete these prints to enhance system performance. It has aided developers in discovering that over 50% of the log print statements in the system are unnecessary. As a result, the performance of Apache IoTDB has improved by approximately 8%.